### v1

# FlashBack : Efficient Retrieval-Augmented Language Modeling for Long Context Inference

Abstract
Retrieval-Augmented Language Modeling (RALM) by integrating large language models (LLM) with relevant documents from an external corpus is a proven method for enabling the LLM to generate information beyond the scope of its pre-training corpus. Previous work using utilizing retrieved content by simply prepending retrieved contents to the input poses a high runtime issue, which degrades the inference efficiency of the LLMs because they fail to use the Key-Value (KV) cache efficiently. In this paper, we propose FlashBack, a modular RALM designed to improve the inference efficiency of RALM with appending context pattern while maintaining decent performance after specific fine-tuning without heavily destruct the knowledge integrity of the LLM. FlashBack appends retrieved documents at the end of the context for efficiently utilizing the KV cache instead of prepending them. Our experiment shows that the inference speed of FlashBack is up to 4×4$\times$4 × faster than the prepending method on a 7B LLM (Llama 2). Via bypassing unnecessary re-computation, it demonstrates an advancement by achieving significantly faster inference speed, and this heightened efficiency will substantially reduce inferential cost. Our code will be publicly available.
1 Introduction
Figure 1: A series of inference tests with simulated input and retrieved content on OPT and Llama 2 for comparing Appending Pattern (ours) with Prepending Pattern. (Maximum sequence length is denoted under the model name)
Large language models (LLMs) based on the Transformer architecture (Vaswani et al., 2023) such as GPT, Llama and OPT, etc (Brown et al., 2020; Touvron et al., 2023; Zhang et al., 2022) require enormous computational resources to keep their knowledge updated (Meng et al., 2023). Thus, Retrieval-Augmented Language Modeling (RALM) has emerged as a popular approach, enabling content generation that leverages external corpora to extend beyond the knowledge inherent in the model’s parameters, thereby reducing the computational cost of capturing up-to-date knowledge. The effectiveness of using RALM to enhance LLMs’ performance has been proven by previous studies(Guu et al., 2020; Izacard et al., 2022; Lewis et al., 2021; Wang et al., 2023b).
Previous studies build RALM through the pre-training of LLM(Borgeaud et al., 2022), or engage in the continual pre-training of existing RALMs (Izacard et al., 2022; Wang et al., 2023a). Recent works utilizing retrieved content by simply prepending retrieved contents to the input without updating the language models exhibit the impressive potential of adapting off-the-shelf LLMs to the retrieved content (Ram et al., 2023; Shi et al., 2023).
However, these works are introduced with limitations. First, the off-the-shelf LLMs are not inherently trained to incorporate retrieved content, and extensive pre-training of LLMs for building RALM incurs high computational costs (Lin et al., 2023). Second, although in-context methods have been effectively applied on off-the-shelf LLMs (Ram et al., 2023; Shi et al., 2023), recent research indicates that the bottleneck of these methods is redundancy and inefficiency (Asai et al., 2024) and our work verified that prepending context pattern they used poses high computational cost in long context scenario, which tremendously increases the inference time. To be specific, retrieved content changes with every retrieval stride tokens, and the computed key-value (KV) cache of the entire context is discarded, requiring re-computation for each change. Text generation process with long context can substantially increase the inference runtime, and the attention KV cache grows even larger on LLM with hundreds of billions of parameters(Pope et al., 2022). Under these circumstances, the re-computation of the KV cache becomes a computationally intensive task, entailing significant costs for In-Context RALM(Ram et al., 2023).
In this paper, we propose FlashBack to overcome these issues we mentioned above, a method that utilize off-the-shelf LLMs efficiently without further pre-training of LLMs as shown in figure 3, to address the inefficiency issue by breaking the dependent of the input to retrieved content in the context. To be specific, instead of prepending the retrieved content directly to the input, we propose a new context pattern, Appending Context Pattern, in which the retrieved content is appended at the end of the input (see figure 2). In our method, the KV cache of each input is static during inference, and the re-computation of the KV cache takes place only in the part of the retrieved content. Although there are advantages to applying this new context pattern, it also presents an issue that degrading the performance of the RALM. Our appending pattern breaks the semantic coherence of the context, which leads to a loss in perplexity when applying it to existing RALM. To solve this issue, inspired by (Ren et al., 2023), FlashBack introduces tunable Marking Token to adapt to our appending context pattern for restoring performance, which is different from pre-training.
Notably, we use LoRA (Hu et al., 2021) to fine-tune the Marking Token while keeping both the retriever and the LLM frozen. Therefore, FlashBack is a plug-and-play approach that can cooperate with other retrieval-oriented methods aiming for efficient inference and context-model alignment. When using an LLM with 7 billion parameters (7B) as a reader, user can run the whole token-tuning process with a single 24GB-RAM GPU in BF16 precision.
Our main contributions can be listed as follows:
  • We find the Appending Context Pattern can improve the inference efficiency of In-Context RALMs. (Up to 4×4$\times$4 × faster in 7B Llama 2 as shown in figure 1)
  • We exploit Marking Token that can be used for adapting the LLMs to unseen context pattern by slight fine-tuning with LoRA. And this alignment method with LoRA can keep the LLM intact and avoid heavily damaging the knowledge integrity of the LLM.
  • Our modeling, the FlashBack, is orthogonal to other methods using off-the-shelf LLM. It has considerable potential for enhancement combined with other methods in future works.
Figure 2: An illustration of the Appending Context Pattern(ours) comparing with Prepending Context Pattern.
Figure 3: An overview of the FlashBack pipelines. The left side of the diagram is the fine-tuning phase of our model with Marking Token and LoRA module. The inference procedure is on the right.
2 Background and Related Work
2.1 Retrieval-Augmented Language Models
$k$NN-LM is a pioneering method that substantiated its capability as RALM, and it suggests that learning similarity functions between contexts may be a better solution than predicting the next word (Khandelwal et al., 2020). However, a recent investigation (Wang et al., 2023c) has revealed that the perplexity of $k$NN-LM exhibits improvement for a limited set of tokens but exacerbates predictions for the majority of tokens, particularly when generating lengthy sequences. As a result, this characteristic adversely affects the overall quality of text generation when compared to GPT-2 (Wang et al., 2023c). RALM also can be based on models using encoder-decoder structure(Huang et al., 2023; Lewis et al., 2021) and Atlas (Izacard et al., 2022) building upon the T5 language model (Raffel et al., 2023) stands out as a state-of-art RALM. Other RALMs (Lin et al., 2023; Ram et al., 2023; Shi et al., 2023; Xu et al., 2023) studied for utilizing decoder-only language models such as GPT, Llama and OPT, etc(Brown et al., 2020; Touvron et al., 2023; Zhang et al., 2022) or building their own autoregressive RALM (Borgeaud et al., 2022).
2.2 Retrieve-Read RALM
Previous works (Borgeaud et al., 2022; Lin et al., 2023; Ram et al., 2023; Shi et al., 2023) have created distinct modules for document selection and document reading. In the recent RALM framework, particularly for those employing LLMs, the imperative is to align retrieved documents with the specific requirements of the LLMs (Gao et al., 2024). AAR, REPLUG, and UPRISE are fine-tuning retrievers with frozen LLMs to achieve the alignment (Cheng et al., 2023; Shi et al., 2023; Yu et al., 2023). RETRO uses frozen retrievers to build their own RALM (Borgeaud et al., 2022). In-Context RALM uses a frozen retriever for document selection and a frozen LLM for document reading without undergoing additional training for either the LLM or the retriever (Ram et al., 2023). RAVEN (Huang et al., 2023) continually pre-train ATLAS (Izacard et al., 2022) for improving its in-context capabilities, which aligns the model to specific prompting strategy, and they also proposed Fusion-in-Context Learning for incorporating more in-context examples during its inference process.
Hence, our modeling is close to the genre of RALM that uses decoder-only LLMs (Auto-regressive Models) and Retrieval-Read RALM that uses distinct modules for document selection and document reading (Borgeaud et al., 2022; Ram et al., 2023; Lin et al., 2023; Shi et al., 2023). Also, our implementation on input context is classified as Input augmentation by recent study(Asai et al., 2024). After all, This type of RALM has recently been defined as Modular RAG (Retrieval-Based-Generation), and it is increasingly becoming the dominant norm in the RAG domain (Gao et al., 2024).
Figure 4: Comparison of prepending and appending context patterns in interval of change of retrieved content, arrow represents that generated tokens in past round will become a part of input in the next round. In the prepending context pattern,the key-value of all tokens in past round should be re-computed. By contrast, in appending context pattern, the key-value of input in past round can be reused in the next round.
3 Methodology
3.1 RALM with In-Context-Learning
In the In-Context RALM framework (Ram et al., 2023), an external corpus $\mathcal{C}$ is provided to the retriever. The retriever is responsible for determining the specific content within the corpus to be utilized as the retrieved content, and then the retrieved content can be concatenated with input to form a context that conditions the predictions of the LLM. Given a sequence of tokens $x_{1},...,x_{n}$ with length $n$, The probability of token sequence is represented by
$p(x_{1},...,x_{n})=\prod_{i=1}^{n}p_{\theta}(x_{i}|[\mathcal{R}_{\mathcal{C}}(x_{<i});x_{<i}]),$ (1)
where $\theta$ is parameters of an auto-regressive model, $\mathcal{R}_{\mathcal{C}}(x_{<i})$ is retrieved content from corpus $\mathcal{C}$ using prefix $x_{<i}$ and $[\mathcal{R}_{\mathcal{C}}(x_{<i});x_{<i}]$ denotes the concatenation of strings $\mathcal{R}_{\mathcal{C}}(x_{<i})$ and $x_{<i}$.
For the trade-off between runtime cost and model performance, retrieval stride $s$ and query length $\ell$ are proposed for less frequently performing the retriever call (at least, every s > 1 token). Therefore, the retrieval would be performed for every $s$ tokens, using the last $\ell$ tokens as designated retrieved content for input.
The probability of the token sequence is represented by
$p(x_{1},...,x_{n})=\prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)});x_{<s\cdot j+i}]),$ (2)
where $n_{s}=n/s$ is the number of retrieval strides, $q_{j}^{(s,\ell)}=x_{s\cdot j-\ell+1},...,x_{s\cdot j}$ is the retriever input.
Our work followed this mechanism of retrieval besides using different context patterns because this mechanism uses distinct retrieve modules and readers (LLMs) so that they can be optimized independently.
3.2 Context pattern
In the decoder-only transformer-based models, the computation of attention modules is related to the query of the current token and the key-value representations of preceding tokens. (Vaswani et al., 2023). During inference, the key-value of the current token would be saved in the KV cache, eliminating the requirement in the subsequent step. We find that prepending retrieved content to the input has been a prevalent in previous methods(Ram et al., 2023; Shi et al., 2023). However, the key-value computed for previous tokens becomes obsolete, and a re-computation of new key-value is required for each retrieval step since the retrieved content varies with each retrieval.
To avoid the extensive cost of re-computation, we remove the auto-regressive dependency between retrieved content and input in context by using an appending context pattern. Specifically, we append retrieved content to input (at the end of the input) as shown in Figure 4.
Then, the probability of the token sequence is represented by
$p(x_{1},...,x_{n})=\prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[x_{<s\cdot j+i};\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})]).$ (3)
In this case, for each input, the change of retrieved content $\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})$ would not require a re-computation of new value therefore preserving the computed value of previous tokens in KV cache $x_{<s\cdot j+i}$. (Algorithm pseudo code in Appendix A.)
Figure 5: Time for OPT models generating a sequence of length 2048 tokens (including 128 tokens for retrieved content), given a input of 512 tokens, with retrieval stride of 16.
3.3 FLOPs analysis of context pattern
[…]
$C_{0}=2l\sum_{i=1}^{T/s}2bish^{2}=\frac{2T(T+s)bh^{2}l}{s}.$ (4)
It increases quadratically as a function of the maximum sequence length $T$, and thus, it will increase dramatically when the maximum sequence length $T$ is at a large value.
3.4 Marking Token and Fine-tuning Choice
Since LLMs are not aligned explicitly to our appending pattern, we use Marking Token and LoRA techniques to adapt them to the appending pattern while keeping origin model weights frozen so that the alignment is achieved without modifying the inherent ability of the LLMs. Marking Tokens are represented as two special prompt tokens denoting <MARK_L>,<MARK_R> which are added into the vocabulary of the model, and they are used for marking the boundary of the retrieved content in context (see Figure 3). Then, we fine-tune the model using Low-Rank Adaption (LoRA) (Hu et al., 2021) with Marking Tokens applied for adapting the appending pattern. LoRA is a parameter-efficient fine-tuning (PEFT) for fine-tuning a subset of parameters in the pre-trained model while targeting for obtaining comparable performance to full model fine-tuning. Recent study suggest that forgetting is inevitable in fine-tuning but PEFT method can facilitate less forgetting during fine-tuning process which does not heavily damage the knowledge integrity of the pre-trained LLM (Kalajdzievski, 2024).In FlashBack, we opt for LoRA with the aim of demonstrating how fine-tuning a relatively minimal set of model weights enables our model to adapt to the appending context pattern while keeping pre-training model weights frozen. Our experiments show that FlashBack has significantly boosted the inference speed, and it can still maintain decent or better perplexity (PPL) after fine-tuning. During fine-tuning, all parameters are frozen except for the embedding of <MARK_L>, <MARL_R>, and LoRA modules applied to attention layers.
3.5 FLOPs analysis of Marking Token and LoRA
If the rank of weight matrices of LoRA is $r$, then the projection matrix dimensions are $[h,r]$. The average length of retrieved content is $d$. We only equip LoRA modules, two projection matrices, for Key and Value matrices. Overall, the FLOPs of LoRA modules in appending context pattern can be formulated as: [^1] We only consider multiplication operations and omit some addition operations.
$\begin{split}C_{1}&=2l(4bThr+bTh)+2l\sum_{i=1}^{T/s}\left[4b(d+s)hr+b(d+s)h\right]\\ &=\frac{2l(4r+1)bhT(d+2s)}{s}\end{split}$ (5)
Combining equation 4 and 5, the decreament of FLOPs when using appending context pattern with LoRA is:
$C_{\text{decreament}}=C_{0}-C_{1}=\frac{2lTbh\left[(T+s)h-(4r+1)(d+2s)\right]}{s}$ (6)
3.6 Retriever
Our experiment used a sparse model, the BM25 (Robertson and Zaragoza, 2009), for demonstrating our idea. In practice, any other retrievers can also be used for the document selection process, and FlashBack intrinsically supports switching to other retrievers in a plug-and-play manner.
Model A./P. LoRA M.T. Wikitext-2 Arxiv Freelaw Stackexchange
GPT2-774M N. 19.11 10.69 11.29 9.74
P. 16.71 10.58 11.08 9.59
P. 11.74 8.56 8.03 6.45
P. 11.69 8.52 8.03 6.43
A. 56.15 42.44 49.25 39.61
A. 14.40 16.83 16.35 11.99
A. 12.13 8.95 8.47 6.78
GPT2-1.5B N. 17.32 9.87 10.61 9.03
P. 15.36 9.81 10.49 9.01
P. 10.84 8.07 7.45 6.01
P. 10.83 8.04 7.43 5.98
A. 53.30 41.72 49.25 40.25
A. 13.75 16.48 15.20 11.50
A. 11.32 8.57 8.01 6.35
OPT-1.3B N. 16.76 9.64 8.58 7.78
P. 15.02 9.59 8.35 7.63
P. 10.23 8.33 7.47 6.87
P. 10.22 8.31 7.45 6.87
A. 81.57 51.94 53.73 42.99
A. 12.65 11.38 10.11 9.13
A. 10.49 8.72 7.85 7.17
OPT-2.7B N. 14.50 8.72 7.74 6.99
P. 13.14 8.68 7.56 6.88
P. 9.23 7.69 6.80 6.23
P. 9.23 7.68 6.78 6.22
A. 76.41 50.68 51.78 42.48
A. 11.58 10.88 9.31 8.51
A. 9.52 8.08 7.19 6.53
OPT-6.7B N. 12.30 7.74 6.94 6.22
P. 11.20 7.73 6.83 6.15
P. 8.23 6.98 6.19 5.58
P. 8.24 6.99 6.18 5.58
A. 68.31 46.53 48.33 40.25
A. 10.54 10.92 9.27 8.04
A. 8.59 7.43 6.64 5.94
Table 1: Perplexity (the lower the better) of GPT2 and OPT models on a part of the validation set of WikiText-2 dataset and the test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refer to the Marking Tokens. And N. refers to none retrieval.
4 Experiments
4.1 Run-time Improvement
Setup
We test OPT, GPT-2, and Llama2 in different model sizes and employ simulated retrieved content and input in inference speed tests. Our experiment encompassed tests under varying lengths of input. Additionally, it involved examining the inference time under two distinct context patterns.
Figure 6: Time for Llama2-7B generating sequences with different length, given a input of 512 tokens, with retrieval stride of 16.
Results
In Figure 5, we compare the appending pattern and prepending pattern in OPT models. We scale the model size from 125M to 6.7B. It is obvious that the acceleration of inference time is more effective in large models that have more layers and larger hidden sizes. In figure 6, we scale the max sequence length from 1088 to 3968 and observe significant improvement, as Llama 2 has a larger max sequence length, which is 4096 tokens. (Note: The test results for inference speed may vary depending on the hardware used; our runtime tests are conducted on a single A100-80GB GPU)
4.2 Language Modeling
Setup
We test FlashBack on Wikitext-2 (Merity et al., 2016) dataset, and three dataset: Arxiv, Freelaw and Stackexchange from the Pile (Gao et al., 2020). We report token-level average perplexity as the evaluation metric (excluding <MARK_L>,<MARK_R>). Our experiments use models of OPT (1.3B-6.7B) and GPT-2 (124M-1.5B). We compare FlashBack with the appending context pattern without fine-tuning and also the prepending context pattern. (For smaller models that are under 774M parameters, the results are in Appendix Table 6)
Results
Table 1 shows the results of different setup on the dev set. We can see that the appending context pattern applied directly on LLMs through the in-context method generates higher perplexity compared with the model using the prepending context pattern. Our result on fine-tuning with appending context patterns by using LoRA demonstrates that the model can adapt to such patterns after fine-tuning. Notably, with both the Marking Token and LoRA applied, the difference between perplexity results of two different context patterns is increasingly mitigated as the test model size grows larger.
4.3 Language Modeling with more retrieved content
Inspired by REPLUG(Shi et al., 2023), we append each retrieved content separately to the input and ensemble respective output probabilities.
Setup
We use the same setup as section 4.2, except varying the number of retrieved content.
Perplexity
Model # Of Retrieved Content
1 2 4 8
GPT2-124M 25.00 24.06 24.65 25.72
GPT2-355M 16.05 15.58 15.91 16.50
GPT2-774M 13.28 12.94 13.21 13.67
GPT2-1.5B 12.49 12.19 12.43 12.83
OPT-125M 23.44 22.53 23.09 23.97
OPT-350M 16.89 16.31 16.67 17.29
OPT-1.3B 11.45 11.10 11.32 11.67
OPT-2.7B 10.48 10.14 10.32 10.63
OPT-6.7B 9.34 9.05 9.20 9.34
Table 2: Perplexity (the lower the better) of GPT models and OPT models on the validation set of WikiText-2 dataset, varying the number of retrieved content, by using FlashBack.
Results
From the test results of varying the number of retrieved contents, the perplexity of all models we tested slightly worsens as the number of retrieved contents increases. We find the best perplexity comes with setting the number of retrieved contents to 2. We speculate the reasoning of this phenomena is that the second retrieved document is likely to be highly relevant with the first retrieved document. However, this experiment is only for demonstrating that FlashBack can be used for multi-documents retrieval. Further improvement on multi-documents retrieval is not our focus in this paper and it can be a direction in later work.
4.4 Implementation Details
Run-time Improvement
We randomly sample from a discrete uniform distribution $U\{500,1000\}$ to get simulated token sequences to be the pseudo input and pseudo retrieved content. We maintain the length of simulated retrieved content to a fixed value of 128, then test models with different maximum sequence lengths. (Setup details are shown in Appendix Table 3). For The retrieval stride, all models are set to 16. All experiments are conducted by using a single GPU.
Language Modeling
For every $s$ tokens in the train set (refer to target), we sample from a discrete uniform distribution $U\{T/2,T-s\}$ to obtain the start position of the target, where $s$ is retrieval stride and $T$ is the maximum sequence length. For Wikitext-2, Wikitext-103, a superset of Wikitext-2, is used as a retrieval corpus. To avoid data leakage that target outputs are included in retrieved contents, we filter out retrieved content with titles that match Wikitext-2 during data construction. Owning to the much larger size of the Pile dataset, for Arxiv, Freelaw and Stackexchange, the validation split and 80% of the test split are mixtured as a retrieval corpus. And the rest of 20% of the test split are divided into two part equally as the training and the test dataset, respectively. We employ the BM25 retrieval and configure the query length to 16, indicating the utilization of the preceding 32 tokens from the target as input for the retriever. And we use the next token prediction objective for language modeling experiments. The loss is computed on target tokens in context.
Training Details
The trainable parameters for all models include two token embedding of <MARK_L> and <MARK_R>, and LoRA modules applied to all attention layers. The rank of weight matrices of LoRA is set to 16. The batch size is set to 16. The learning rates for GPT-2 models (124M-1.5B) are specified as [4e-4, 2e-4, 1e-4, 5e-5]. Correspondingly, the learning rates for OPT models (1.3B-6.7B) are specified as [1e-4, 5e-5, 5e-5]. We train for 5K additional steps, with 10% warm-up steps and a cosine learning rate schedule. We use the bf16 format for the OPT-6.7B model and set the maximum sequence length to 512 for running on a single 24GB RAM GPU.
Our code uses the Transformers library (Wolf et al., 2020) and BM25 retrieval from the Pyserini library (Lin et al., 2021). All fine-tuning experiments are conducted by using 4$\times$24GB RAM GPUs.
5 Conclusion
Prepending retrieved contents to the input without modification of model architecture has been a practicable method for LM to use external knowledge. However, the prepending context pattern hinders RALM from generalizing efficiently in considerable context length.
This paper presented the FlashBack method, enabling In-Context RALM with faster inference speed while keeping competitive model performance. We proved that the appending context pattern is more efficient than the prepending counterpart. Besides, we demonstrated that LLMs can adapt to new context patterns using Marking Tokens with adaption fine-tuning (LoRA) and its performance in generation quality is likely equivalent to the LLM using prepending counterpart as model grows larger.
6 Limitations
In our experiments, the retrieval stride is a fixed value in each test, but it can be a dynamic variable in real applications, so there might be extra improvements. In run-time comparison experiments, due to constrained computational resources, we used simulated (pseudo) data for testing; therefore, we have results for comparing inference speed. Also, It is computationally intensive to fine-tune an LLM over 7B parameters, so the performance of FlashBack in extra large models (over 7B) might require further demonstration.
Frozen LM with LoRA modules explicitly makes it easy to switch modes between LLM and RALM. By just loading LoRA modules, a frozen LLM can use the external corpus and vice versa. In some situations, knowledge stored in model parameters might be superior so that it is possible for RALM to evaluate the quality of the retrieved contents and dynamically determine whether the external corpus should be used in generating the output. Furthermore, in FlashBack, we did not scale the number of retrieved contents to a large value. It is practicable to generalize appending context patterns to more retrieved contents (for example, using the framework of REPLUG(Shi et al., 2023)), which can be one of the future work to do.
These we mentioned above can be future research for RALM.
Appendix A FlashBack Algorithm
Algorithm 1 In-Context RALM with FlashBack
Given retrieve stride $s$, retrieved query length $\ell$, and maximum sequence length $T$.
Given auto-regressive model $p_{\theta}$, retriever $q$, and initial prompt sequence $x_{0},...,x_{t}$.
Initialise: $n\leftarrow t$, retrieved evidence $e\leftarrow\text{None}$, Key-Value Cache $K\leftarrow\text{None}$.
while $n\leq T$ do
     The retriever is called for every stride $s$ tokens.
     if $(n-t\mod s)=0$ then
         $e\leftarrow q(x_{n-\ell,..,n})$triangleright Call retriever
         $l_{e}\leftarrow\text{length of }e$
         Merge past generated tokens to the context, using truncated KV cache:
         $p_{n},K_{1,...,l_{e}+n-1}\leftarrow p_{\theta}(x|[x_{1,...,\tilde{n}-1};x_{\tilde{n},...,n-1};e];K_{1,...,\tilde{n}-1})$
         $\tilde{n}\leftarrow n$triangleright Save last retriever call position
     else
         $p_{n},K_{1,..,l_{e}+n-1}\leftarrow p_{\theta}(x|[x_{1,..,\tilde{n}-1};e;x_{\tilde{n},...,n-1}];K_{1,..,{l_{e}+n-2}})$
     end if
     $x_{n}\sim p_{n}(x)$
     $n\leftarrow n+1$
end while
Appendix B Setup and Experiments
Models Length of input Length limitation
GPT2 256 1024
OPT 512 2048
Llama2 512 4096
Table 3: The length of input and the length limitation we used in run-time improvement experiments.
B.1 Ablation Study
Marking Token
We have already tested the model using an appending context pattern that is solely fine-tuned with LoRA without Marking Token to compare with FlashBack. In Table 1, test results of models with Marking Tokens and LoRA fine-tuning have lower PPL, which means fine-tuning the Marking Token could further boost the performance of the RALM and, therefore, meet our expectations. Further more, we also tested whether Marking Token and LoRA module add additional inference costs to our modeling. Our results on figure 5 show that FlashBack is slightly slower than model without these modules but the difference is tiny enough to be ignored compared to the overall increase of the inference speedup.
Context Pattern
Retrieval Stride
Allocating the number of retrieval strides for each is a design choice that is related to model performance in both perplexity and inference time. As we mentioned above in section 3.1, In-Context RALM [Ram et al., 2023] choose to use a cherry-picked retrieval stride to balance the performance trade-off. They choose to use a relatively small stride value to have better generation quality (lower perplexity) without adding extensive inference cost. However, our experiments (see table 4 and 5 in appendix) on varying retrieval strides on FlashBack implied a different opinion. We find using 32 or 64 as retrieval stride does not degrade the model’s performance. When testing with a larger model such as OPT-6.7B, the test with a higher retrieval stride has the lowest perplexity compared to the test using same model with lower retrieval stride. Therefore, the trade-off between perplexity and inference time may not be a particularly evident phenomenon for some cases. As a consequence, we made a bold speculation that it might be an advantage for FlashBack for increasing the retrieval stride without intensively degrading the model’s generation quality in perplexity and still maintaining high-speed inference in long context scenarios. And this phenomenon may requires further demonstration in the future.
Perplexity
Model Retrieval Stride
16 32 64
GPT2-124M 28.11 23.64 21.80
GPT2-355M 16.46 15.66 15.80
GPT2-774M 13.46 13.12 13.24
GPT2-1.5B 12.59 12.28 12.24
OPT-125M 24.73 22.42 21.26
OPT-350M 17.29 16.71 16.71
OPT-1.3B 11.51 11.29 11.31
OPT-2.7B 10.70 10.38 10.23
OPT-6.7B 9.40 9.10 9.09
Table 4: Perplexity (the lower the better) of GPT models and OPT models on the validation set of WikiText-2 dataset, varying retrieval stride, by using FlashBack.
Time (Second)
Model Retrieval Stride
16 32 64
GPT2-124M 6.16 5.67 5.89
GPT2-355M 10.68 10.37 10.35
GPT2-774M 16.08 15.97 15.77
GPT2-1.5B 21.71 21.23 20.26
OPT-125M 10.72 10.09 10.10
OPT-350M 20.01 19.9 19.69
OPT-1.3B 22.13 20.72 19.79
OPT-2.7B 30.71 29.10 28.17
OPT-6.7B 50.25 43.79 41.06
Table 5: Time for GPT models and OPT models generating sequences with respective maximum context length, varying retrieval stride.
B.2 Perplexity of GPT model in small model size
Model A./P. LoRA M.T. Wikitext-2 Arxiv Freelaw Stackexchange
GPT2-124M N. 30.69 15.19 16.65 14.76
P. 26.00 14.98 16.00 14.43
P. 17.46 11.46 11.55 9.02
P. 17.41 11.44 11.53 9.00
A. 72.29 46.59 56.32 50.71
A. 25.56 22.35 24.31 17.41
A. 19.14 14.21 15.01 10.25
GPT2-355M N. 22.13 11.79 12.01 10.07
P. 19.26 11.68 11.63 9.91
P. 13.33 9.46 9.00 7.13
P. 13.33 9.44 8.97 7.13
A. 59.50 42.33 47.23 37.00
A. 17.05 18.60 18.54 13.18
A. 14.37 10.51 9.94 7.81
Table 6: Perplexity (the lower the better) of GPT2 models on a part of the validation set of WikiText-2 dataset and the test set of Arxiv, Freelaw and Stackexchange datasets. The retrieval stride is set to 16. A. and P. refers to the appending and prepending context pattern respectively. M.T. refers to the Marking Tokens. And N. refers to none retrieval.